OpenStudy (anonymous):

Please help! It would be greatly appreciated. How many solutions does this system have? y - 5x = -6 3y - 15x = -12

OpenStudy (mathstudent55):

Multiply the first eq by -3 and add to the second equation.

OpenStudy (mathstudent55):

-3y +15x = 18 3y - 15x = -12 --------------- 0 + 0 = 6 Since the addition is 0 = 6, and that's a false statement, that means there is no solution.